The problem is this:
(1-2x)(3x+1)-(x-1)^2
What I have done is this:
(x-1)^2 = x^2 -2x +1
And using FOIL, (1-2x)(3x+1) =
3x +1 -6x^2 -2x
So, simplified, the equation reads: -x -7x^2 +2 or -7x^2 -x +2
But the multiple choice answers are:
a. x(3-7x)
b. -7x^2 -x
c. -7x2 -5x
d. -7x^2 +x
e. None of these
Am I right? Is it e. none of these? Please help!
: The problem is this:
(1-2x)(3x+1)-(x-1)^2
What I have done is this:
(x-1)^2 = x^2 -2x +1
And using FOIL, (1-2x)(3x+1) =
3x +1 -6x^2 -2x
So, simplified, the equation reads: -x -7x^2 +2 or -7x^2 -x +2
But the multiple choice answers are:
a. x(3-7x)
b. -7x^2 -x
c. -7x2 -5x
d. -7x^2 +x
e. None of these
Am I right? Is it e. none of these? Please help!

Answer by nerdybill(1280)

(1-2x)(3x+1)-(x-1)^2
What I have done is this:
(x-1)^2 = x^2 -2x +1
And using FOIL, (1-2x)(3x+1) =
3x +1 -6x^2 -2x
So far, so good!!
.
Now, rewriting:
(1-2x)(3x+1)-(x-1)^2
We get:
(3x +1 -6x^2 -2x) - (x^2 -2x +1)
3x +1 -6x^2 -2x - x^2 + 2x - 1
3x -6x^2 -2x - x^2 + 2x
3x -6x^2- x^2
3x -7x^2
factoring out an 'x':
x(3-7x) (Which is answer 'a')